OpenStudy (anonymous):

A dive bomber has a velocity of 255 m/s at an angle θ below the horizontal. When the altitude of the aircraft is 2.15 km, it releases a bomb, which subsequently hits a target on the ground. The magnitude of the displacement from the point of release of the bomb to the target is 3.35 km. Find the angle θ.

OpenStudy (anonymous):

If you do not consider the action of gravity, you will have a triangle: 

OpenStudy (anonymous):

Pretty sure i need to consider gravity as 9.8 m/s^2

OpenStudy (anonymous):

now, tan(theta) = 3.35/2.15 theta = atan(3.35/2.15) theta = 1.00021rad = 57.31º

OpenStudy (anonymous):

Considering the gravity, we must first find the velocity components: Vx = V*sin(theta) Vy = V*cos(theta)

OpenStudy (anonymous):

vy = 30.188 vx = 26.242

OpenStudy (anonymous):

i got that far

OpenStudy (anonymous):

how you got these velocities?

OpenStudy (anonymous):

the same equations you jsut posted..

OpenStudy (anonymous):

that is not correct because 255 is not equal to sqrt(30.188^2 + 26.242^2)

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

L = Vx*t L = V*sin(theta)*t, now find t in terms of theta

OpenStudy (anonymous):

H = Vy*t - 0.5*9.8*t^2 H = V*cos(theta)*t - 0.5*9.8*t^2

OpenStudy (anonymous):

solve for t, plug it back and solve for theta

OpenStudy (anonymous):

what is vx?

OpenStudy (anonymous):

is the x component of teh velocity